OpenStudy (anonymous):

Simplify a - {b - [c - (d - e) - f] - g}. a)a + b + c - d + e - f + g b)a - b + c - d + e - f + g c)a + b + c + d + e - f + g d)a - b + c - d + e + f - g

OpenStudy (anonymous):

original problem: -(d - e) = -d + e, i took each term times -1 next part -[b-d+e-f] = -b+d-e+f] next part -{-b+d-e+f-g} = b-d+e-f+g next part a+b-d+e-f+g the answer